two cars leave the same location at 2.p.m . if one car travels north at the rate of 30 mphs and the other car travels east at the rate of 40 mph. how many miles apart are the two cars at 4 p.m.
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

two cars leave the same location at 2.p.m . if one car travels north at the rate of 30 mphs and the other car travels east at the rate of 40 mph. how many miles apart are the two cars at 4 p.m.
----
Let the cars be "n" and "e":
---
The travel for 2 hrs.
n goes 2*30 = 60 miles
---
e goes 2*40 = 80 miles
----
Use Pythagoras to find the distance between at 4:00PM
-------
d = sqrt[60^2+80^2]
---
d = sqrt[3600 + 6400]
---
d = sqrt(1000)
d = 10*sqrt(10) miles
=======================
Cheers,
Stan H.
=================

Answer by Alan3354(69443)

One car is 60 miles away, the other is 80 miles away.
It's a right triangle.
Use Pythagoras
d = 100 miles